1. Write the equation in slope-intercept form. -1/2x + y = 4
I need help writing this equation in slope-intercept form I'd appreciate the help thanks. If you could explain how you do it that would be great thank-you.

Found 2 solutions by stanbon, Theo:
Answer by stanbon(75887)

Write the equation in slope-intercept form. -1/2x + y = 4
I need help writing this equation in slope-intercept form
----
Form:: y = mx + b
-------
Your Problem::
Solve for "y"::
Add (1/2)y to both sides to get::
y = (1/2)x + 4
------------------
Cheers,
Stan H.
-----------

Answer by Theo(13342)

the slope intercept form of the equation is y = mx + b
m is the slope and b is the y-intercept.

your equation is - (1/2 * x) + y = 4

add (1/2 * x) to both sides of the equation to get:

- (1/2 * x) + (1/2 * x) + y = 4 + (1/2 * x)

the - (1/2 * x) and the + (1/2 * x) on the left side of the equation cancel out and you are left with:

y = 4 + (1/2 * x)

re-order the terms on the right side of the equation in descending order of degree to get:

y = (1/2 * x) + 4

remove the parentheses to get:

y = 1/2 * x + 4

the equation is now in slope intercept form, where:

m = 1/2
b = 4